### v2

@@ II.2 Payoff mechanism
In Appendix  A.4, we show that given some innocuous assumptions, a payoff mechanism can be constructed such that a unique pure strategy Nash equilibrium exists where each player’s dominant strategy is to submit an honest sample from a quantum boson sampler. In other words, we construct a reward and penalty mechanism where it is the player’s unique dominant strategy to behave honestly by committing samples from the boson sampling distribution. This is in contrast to doing nothing, or cheating, i.e. providing samples from some other, presumably easy to calculate, distribution. This situation is described by the inequalities, $u_{i}^{honest}>u_{i}^{nothing}>u_{i}^{cheat}$, representing the net utilities of the $i^{th}$ node under the honest, nothing, and cheating strategies respectively. We set a threshold distance $\epsilon$ between the peak bin probability of individual nodes, $\mu_{i}$, and the net PBP $\mu_{net}$, such that if $f(|\mu_{i}-\mu_{\mathtt{net}}|)<\epsilon$, the nodes are rewarded in proportion to the number of samples they have committed and otherwise are penalized by losing their initial stake. The function $f$ should be monotonic and we can assume it is linear in the argument. The choice of $\epsilon$ decides the tolerance of the rewarding subroutine and can be changed to incorporate multiple error models.
@@
Note that some of these assumptions in this payoff mechanism can be relaxed and still maintain a robust protocol. For example the split reward system can be replaced with a winner takes all block reward, and capital expenditure $k_{fixed}$ for quantum boson samplers can be accounted for by the replacement $k\rightarrow k_{variable}+k_{fixed}/\tau$ in Eq. 37 where $\tau$ is the number of samples the boson sampler is expected to produce before obsolescence.
@@
[…]
@@
The energy cost to run boson samplers is dominated by the cost to cool the detectors since the cost to generate the photons and couple them into the device is negligible. Superconducting single-photon detectors of NbN type with reported efficiencies of $\eta_{d}=0.95$ can operate at $2.1$K [32] which is just below the superfluid transition temperature for Helium. Two-stage Gifford-McMahon cryocoolers can run continuously at a temperature of $2$K with a power consumption of $\sim 1.5$kW [32]. To compare the energy cost of boson-samplers to classical samplers, note that the power consumption of the Tianhe-2 supercomputer is $24$MW [37], and the power consumption of a single core processor at $3.5$GHz is $\sim 100$W. Ultimately, the important metric is the energy cost per sample since it is the accumulation of valid samples that enables a consensus to be reached. As seen from Fig.5, quantum boson-samplers are significantly more energy efficient than classical computers. For example, at $N=25$ photons the quantum boson-sampler costs $6.77\times 10^{-2}$ per sample which is $1563$ times more energy efficient than a single core processor and $29569$ times more efficient than a supercomputer.
@@
While classical devices, such as ASICs, could be developed in the future that would speed up calculations of matrix permanents by some constant factor, any such device is fundamentally going to be limited by the exponential in $N$ slowdown in the sampling rate ($R_{c}$ in Eq. 39). Even as classical computers do speed up, one can increase the number of photons to maintain the same level of quantum advantage. Importantly, this would not require frequent upgrades on the boson sampler since the same device can accommodate a few more input photons as the number of modes was already assumed to be $O(N^{2})$. Furthermore, as the quality of the physical components used for boson-sampling improves, the quantum repetition rates ($R_{q}$ in Eq. 38) will increase, ultimately limited by the single photon source rate.
@@
On the other hand, it is unlikely that much faster “quantum ASIC” devices will be developed for boson sampling. Fock state Boson sampling can be simulated fault tolerantly by universal quantum computers with polynomial overhead. One way to do this is to represent the state space as a truncated Fock space encoded in $M$ qudits of local dimension $N+1$ (or in $M\times\lceil\log{(N+1)}\rceil$ qubits). The input state is a tensor product state of |$\ket{0}$ and |$\ket{1}$ states, the gates of the linear interferometer are two qudit gates which can be simulated in $O(N^{4})$ elementary single and two qudit gates, and the measurement consists of local projectors such that the total simulation scales like $O(N^{4}M^{2})$. Another way to translate boson sampling to quantum circuits performs a mapping to the symmetric space of qudits as described in [38]. Given the algorithmic penalty as well as the gate overheads for error correction, the quantum computer simulation would be slower than a photonic based native boson sampler, except in the limit of very large $N$ where the fault tolerance of the former enables a speedup. However, at that point the entire protocol would be too slow to be of use for consensus anyway.
@@ III Discussion
We have proposed a PoW consensus protocol that natively makes use of the quantum speedup afforded by boson-samplers. The method requires that miners perform full boson-sampling, where samples are post-processed as coarse-grained boson-sampling using a binning strategy only known after samples have been committed to the network. This allows efficient validation but resists pre-computation either classically or quantum mechanically.
@@
The quantum hardware required for the implementation of our protocol has already been experimentally demonstrated at a sufficient scale and is becoming commercially available. While we have focused our analysis primarily on conventional Fock state boson-sampling, the method extends to Gaussian boson-sampling, accommodating faster quantum sampling rates owing to the relative ease with which the required squeezed vacuum input states can be prepared. We have used conservative estimates of the number of samples needed to achieve concensus based and it is plausible that these upper bounds can be tightened. We leave a detailed study of the number of samples required, error tolerances, and performance of Gaussian boson-samplers to future work.
@@ Acknowledgements.
We gratefully acknowledge discussions with Louis Tessler, Simon Devitt, and Peter Turner. GKB and GM received support from a BTQ-funded grant with Macquarie University. GKB and PPR receive support from the Australian Research Council through the Centre of Excellence for Engineered Quantum Systems (project CE170100009). DS is supported by the Australian Research Council (ARC) through the Centre of Excellence for Quantum Computation and Communication Technology (project CE170100012).
@@ A.1.2 Hash functions
The alternative we describe here is to use non universal quantum devices that solve the boson sampling problem for PoW. Unlike the one-way functions described above, classical or quantum verification of boson-sampling is not known to be possible and hence it would appear hopeless as a means to reach concensus. However, by coarse graining the distribution, verification indeed is possible. State-binned boson-sampling (see Sec. I.1.3) was motivated as an attempt to construct a hash function from the boson-sampling problem [10]. This type of function determined by sampling differs from conventional hash functions as it is not in NP, since a classical verifier cannot efficiently verify the output to the hash given the input state. However, binned distributions converge with sufficient samples, enabling verification by quantum verifiers by comparing for consistency in binned distributions. Moreover, using a different kind of course graining by binning output modes, an efficient classical verification can be achieved. We combine both CGBS strategies in the full protocol to check that miners are producing valid samples and to provide rewards to successful miners in a manner that is responsive to the combined sampling power of the network.
@@ A.3 Characteristic function for Gaussian Boson Sampling
In order to develop a validation protocol for Gaussian Boson Sampling similar to that for Fock state, we need to calculate the corresponding characteristic function. As before, the characteristic function is defined as,
@@ A.4.6 Components of Costs (Variable $k$) and Cost to Entry
Two comments are worth adding here on the adoption of this new PoW consensus protocol. First, in the early stages, before large scale production and availability of boson samplers, it could be expected that classical miners would dominate. This could be accommodated for by having the reward inequality in Eq. 106 be initially $R>k^{\mathtt{classical}}$ so that the utility of classical players is positive. Then a decision could be made to gradually either (1) increase $k^{\mathtt{classical}}$ (such as increasing the number of photons in the sampling problem and hence the difficulty) or (2) reduce $R$. This will kick classical players out of the protocol as they no longer have positive utility. Second, the conditions on reward and penalty described above assume that the Nash equilibrium is already reached since it is defined by the condition that no unilateral deviation will move the equilibrium. This will not be the case during the initialization stage of the protocol. During the genesis block and several blocks thereafter, additional mechanisms should be placed by trustworthy players, to ensure the initialization reaches this Nash equilibrium. The trustworthy players can then exit the market and the equilibrium will be retained, thus ensuring no “central authority” exists in the protocol.
@@ A.5 Parameters and description
Notation
Description
Comments
BS setup $N$
Total number of input photons
$M$
Total number of modes
We use $M=O(N^{2})$
$U$
Matrix description of linear optical circuit
$U\in\mathbb{C}^{M\times M}$
State-binned BS $d^{(\mathtt{sb})}$
@@
Number of bins in the state Hilbert space
$d^{(\mathtt{sb})}\lesssim 0.1/2\epsilon^{0.8}$
$\mu_{\mathtt{true}}$
Peak bin probability (PBP) of state-binned BS
$\mu_{\mathtt{net}}$
Estimated PBP from all validated samples on the network
$\epsilon$
Desired accuracy of the estimated PBP wrt. $\mu_{\mathtt{true}}$
$N^{(\mathtt{sb})}_{tot}$
@@
The number of samples needed across the network such that $|\mu_{\mathtt{net}}-\mu_{\mathtt{true}}|\leq\epsilon$ with high confidence
$N^{(\mathtt{sb})}_{\rm tot}=1.8\times 10^{5}{d^{(\mathtt{sb})}}^{7/2}$
$\gamma$
$100(1-\gamma)\%$ is the confidence interval for $\mu_{\mathtt{true}}$
We assume $\gamma=10^{-4}$
Mode-binned BS $d^{(\mathtt{mb})}$
Number of bins in the mode Hilbert space
$P^{({\mathtt{m}b})}[i]$
Mode-binned probability distribution for miner $i$
$m$
@@
Number of random samples needed to estimate matrix permanent up to an $\delta$ additive estimate with probability at least $p$, using Gurvit’s algorithm
$m=\frac{2}{\delta^{2}}\ln(2/(1-p))$
$\beta$
@@
Accuracy parameter used to invalidate a miner’s mode binned distribution based on total variation distance from an estimated mode binned distribution $\widehat{P^{({\mathtt{mb}})}}$ calculated using Gurvit’s algorithm
$\mathcal{D}^{({\rm tv})}(\widehat{P^{({\mathtt{mb}})}},P^{({\mathtt{mb}})}[i])\geq 2\beta$
$N^{(\mathtt{mb})}_{tot}$
@@
The number of samples sufficient for a quantum boson sampler to pass the validation test
$N^{(\mathtt{mb})}_{{\rm tot}}=2^{14}\frac{\sqrt{\binom{N+d^{(\mathtt{mb})}-1}{N}}}{\beta^{2}}$
$m_{j}[i]$
Number of photons in the $j^{th}$ bin for the $i^{th}$ sample set $s_{i}$
$s_{i}^{(v)}$
@@
Set of samples committed by the $i^{th}$ node in the blockchain network.
The $v$ superscript, if any, denotes the samples are validated
$W^{(v)}$
The combined set of samples committed by all the nodes in the blockchain network
$W^{(v)}=\bigcup_{i}s^{(v)}_{i}$.
Experimental resources $\eta_{f}$
@@
Combined parameter for single photon generation, coupling, and detection efficiency
$\eta_{t}$
Transmission probability of a single beam-splitter
$R_{0}$
Single photon repetition rate
$R_{q}$
Repetition rate of a quantum boson sampler
$R_{q}=(\eta_{f}\eta^{M}_{t})^{N}R_{0}/Ne$
Payoff mechanism $\mu_{i}$
Peak bin probability over the $i^{th}$ verified sample set $s_{i}^{v}$
$t_{i}$
Time when the $i^{th}$ user commits their samples to the network
$p_{i}^{\mathtt{honest}}$
@@
Probability of an honest user (either classical or quantum sampler) to pass the mode-binned BS test
$p_{i}^{\mathtt{cheat}}$
@@
Probability of a user providing non-BS samples to pass the mode-binned BS test
$u_{i}^{\mathtt{honest}}$
Utility of an honest node in the network
$u_{i}^{\mathtt{cheat}}$
Utility of a cheating node in the network
$u_{i}^{\mathtt{nothing}}$
Utility of doing nothing, i.e. not committing any samples
$k^{\mathtt{classical}}$
Cost per honest sample using classical algorithms
$k^{\mathtt{quantum}}\equiv k$
@@
Cost per honest sample using actual BS implementation
$C_{i}$
The cost incurred by the $i^{th}$ node in committing $|s_{i}|$ samples
$P_{i}$
Penalty implemented on the $i^{th}$ node in the network
$R_{i}$
Reward awarded to the $i^{th}$ node in the network

### v3

@@ II.2 Payoff mechanism
In Appendix A.4, we show that given some innocuous assumptions, a payoff mechanism can be constructed such that a unique pure strategy Nash equilibrium exists where each player’s dominant strategy is to submit an honest sample from a quantum boson sampler. In other words, we construct a reward and penalty mechanism where it is the player’s unique dominant strategy to behave honestly by committing samples from the boson sampling distribution. This is in contrast to doing nothing, or cheating, i.e. providing samples from some other, presumably easy-to-calculate, distribution. This situation is described by the inequalities, $u_{i}^{honest}>u_{i}^{nothing}>u_{i}^{cheat}$, representing the net utilities of the $i^{th}$ node under the honest, nothing, and cheating strategies respectively. We set a threshold distance $\epsilon$ between the peak bin probability of individual nodes, $\mu_{i}$, and the net PBP $\mu_{net}$, such that if $f(|\mu_{i}-\mu_{\mathtt{net}}|)<\epsilon$, the nodes are rewarded in proportion to the number of samples they have committed and otherwise are penalized by losing their initial stake. The function $f$ should be monotonic and we can assume it is linear in the argument. The choice of $\epsilon$ decides the tolerance of the rewarding subroutine and can be changed to incorporate multiple error models.
@@
Note that some of these assumptions in this payoff mechanism can be relaxed and still maintain a robust protocol. For example, the split reward system can be replaced with a winner takes all block reward, and capital expenditure $k_{fixed}$ for quantum boson samplers can be accounted for by the replacement $k\rightarrow k_{variable}+k_{fixed}/\tau$ in Eq. 39 where $\tau$ is the number of samples the boson sampler is expected to produce before obsolescence.
@@
Here $R_{0}$ is the single photon source rate and $R_{0}/N$ is the rate at which $N$ indistinguishable photons are produced, $\eta_{f}$ is a parameter that doesn’t scale with the number of modes and accounts for the preparation and detection efficiencies per photon. It can be written as the product $\eta_{f}=\eta_{g}\eta_{c}\eta_{d}$, where $\eta_{g}$ is the photon generation efficiency, $\eta_{c}$ is the coupling efficiency, and $\eta_{d}$ is the detector efficiency. Finally, $\eta_{t}$ is the beamsplitter transmission probability. Since we are assuming a circuit of depth equal to the number of modes (which is sufficient to produce an arbitrary linear optical transformation), the overall transmission probability per photon through the circuit is $\eta_{t}^{M}$. Finally, the factor of $e$ is an approximation of the probability of obtaining a collision-free event [33]. The experiment of Ref. [34] produced a single photon repetition rate of $R_{0}=76$MHz and the experiment of Ref. [26], reported a transmission probability per photon through a $144\times 144$ optical circuit of $97\%$ implying a per beamsplitter transmission probability of $\eta_{t}=0.97^{1/144}$ as well as an average wavepacket overlap of $99.5\%$, demonstrating high indistinguishability. A photon generation efficiency of $\eta_{g}=0.84$ was reported for quantum dot sources in Ref. [35] and efficiencies of $\eta_{c}=0.9843$ have been demonstrated for coupling single photons from a quantum dot into a photonic crystal waveguide [35]. Finally, single photon detector efficiencies of up to $\eta_{d}=0.98$ have been reported at telecom wavelengths [36]. All these numbers can reasonably be expected to improve as technology advances [37, 38].
@@
The energy cost to run boson samplers is dominated by the cost to cool the detectors and the single-photon sources. Superconducting single-photon detectors of NbN type with reported efficiencies of $\eta_{d}=0.95$ can operate at $2.1$K [36], which is just below the superfluid transition temperature for Helium, and semiconductor quantum dot sources can run at high efficiency and indistinguishability below $10$K [43]. Two-stage Gifford-McMahon cryocoolers can run continuously at a temperature of $2$K with a power consumption of $\sim 1.5$kW [36]. To compare the energy cost of boson-samplers to classical samplers, note that the power consumption of the Tianhe-2 supercomputer is $24$MW [42], and the power consumption of a single core processor at $3.5$GHz is $\sim 100$W. Ultimately, the important metric is the energy cost per sample since it is the accumulation of valid samples that enables a consensus to be reached. As seen from Fig.7, quantum boson-samplers are significantly more energy efficient than classical computers. For example, at $N=25$ photons the quantum boson-sampler costs $6.77\times 10^{-2}$ per sample which is $1563$ times more energy efficient than a single core processor and $29569$ times more efficient than a supercomputer.
@@
While classical devices, such as ASICs, could be developed in the future that would speed up calculations of matrix permanents by some constant factor, any such device is fundamentally going to be limited by the exponential in $N$ slowdown in the sampling rate ($R_{c}$ in Eq. 41). Even as classical computers do speed up, one can increase the number of photons to maintain the same level of quantum advantage. Importantly, this would not require frequent upgrades on the boson sampler since the same device can accommodate a few more input photons as the number of modes was already assumed to be $O(N^{2})$. Furthermore, as the quality of the physical components used for boson sampling improves, the quantum repetition rates ($R_{q}$ in Eq. 40) will increase, ultimately limited by the single photon source rate.
@@
On the other hand, it is unlikely that much faster “quantum ASIC” devices will be developed for boson sampling. Fock state Boson sampling can be simulated fault tolerantly by universal quantum computers with polynomial overhead. One way to do this is to represent the state space as a truncated Fock space encoded in $M$ qudits of local dimension $N+1$ (or in $M\times\lceil\log{(N+1)}\rceil$ qubits). The input state is a tensor product state of |$\ket{0}$ and |$\ket{1}$ states, the gates of the linear interferometer are two qudit gates which can be simulated in $O(N^{4})$ elementary single and two qudit gates, and the measurement consists of local projectors such that the total simulation scales like $O(N^{4}M^{2})$. Another way to translate boson sampling to quantum circuits performs a mapping to the symmetric space of qudits as described in [44]. Given the algorithmic penalty as well as the gate overheads for error correction, the quantum computer simulation would be slower than a photonic-based native boson sampler, except in the limit of very large $N$ where the fault tolerance of the former enables a speedup. However, at that point, the entire protocol would be too slow to be of use for consensus anyway.
@@ III Discussion
We have proposed a PoW consensus protocol that natively makes use of the quantum speedup afforded by boson samplers. The method requires that miners perform full boson sampling, where samples are post-processed as coarse-grained boson sampling using a binning strategy only known after samples have been committed to the network. This allows efficient validation but resists pre-computation either classically or quantum mechanically.
@@
The quantum hardware required for the implementation of our protocol has already been experimentally demonstrated at a sufficient scale and is becoming commercially available. While we have focused our analysis primarily on conventional Fock state boson sampling, the method extends to Gaussian boson sampling, accommodating faster quantum sampling rates owing to the relative ease with which the required squeezed vacuum input states can be prepared. We have used conservative estimates of the number of samples needed to achieve consensus and it is plausible that these upper bounds can be tightened. We leave a detailed study of the number of samples required, error tolerances, and performance of Gaussian boson samplers to future work.
@@ Acknowledgements.
We gratefully acknowledge discussions with Louis Tessler, Simon Devitt, Peter Turner, Jelmer Renema, and Sanaa Sharma. GKB and GM received support from a BTQ-funded grant with Macquarie University. GKB and PPR receive support from the Australian Research Council through the Centre of Excellence for Engineered Quantum Systems (project CE170100009). DS is supported by the Australian Research Council (ARC) through the Centre of Excellence for Quantum Computation and Communication Technology (project CE170100012).
@@ A.1.2 Hash functions
The alternative we describe here is to use non-universal quantum devices that solve the boson sampling problem for PoW. Unlike the one-way functions described above, classical or quantum verification of boson sampling is not known to be possible and hence it would appear hopeless as a means to reach consensus. However, by coarse-graining the distribution, verification indeed is possible. State-binned boson-sampling (see Sec. I.1.3) was motivated as an attempt to construct a hash function from the boson-sampling problem [10]. This type of function determined by sampling differs from conventional hash functions as it is not in NP, since a classical verifier cannot efficiently verify the output to the hash given the input state. However, binned distributions converge with sufficient samples, enabling verification by quantum verifiers by comparing for consistency in binned distributions. Moreover, using a different kind of course graining by binning output modes, an efficient classical verification can be achieved. We combine both CGBS strategies in the full protocol to check that miners are producing valid samples and to provide rewards to successful miners in a manner that is responsive to the combined sampling power of the network.
@@ A.3 Characteristic function for Gaussian Boson Sampling
In order to develop a validation protocol for Gaussian Boson Sampling similar to that for the Fock state, we need to calculate the corresponding characteristic function. As before, the characteristic function is defined as,
@@ A.4.6 Components of Costs (Variable $k$) and Cost to Entry
Two comments are worth adding here on the adoption of this new PoW consensus protocol. First, in the early stages, before large-scale production and availability of boson samplers, it could be expected that classical miners would dominate. This could be accommodated by having the reward inequality in Eq. 98 be initially $R>k^{\mathtt{classical}}$ so that the utility of classical players is positive. Then a decision could be made to gradually either (1) increase $k^{\mathtt{classical}}$ (such as increasing the number of photons in the sampling problem and hence the difficulty) or (2) reduce $R$. This will kick classical players out of the protocol as they no longer have positive utility. Second, the conditions on reward and penalty described above assume that the Nash equilibrium is already reached since it is defined by the condition that no unilateral deviation will move the equilibrium. This will not be the case during the initialization stage of the protocol. During the genesis block and several blocks thereafter, additional mechanisms should be placed by trustworthy players, to ensure the initialization reaches this Nash equilibrium. The trustworthy players can then exit the market and the equilibrium will be retained, thus ensuring no “central authority” exists in the protocol.
@@ A.5 Parameters and description
Notation Description Comments
BS setup $N$ Total number of input photons
$M$ Total number of modes We use $M=O(N^{2})$
$U$ Matrix description of linear optical circuit $U\in\mathbb{C}^{M\times M}$
@@
State-binned BS $d^{(\mathtt{sb})}$ Number of bins in the state Hilbert space $d^{(\mathtt{sb})}\lesssim 0.1/2\epsilon^{0.8}$
$\mu_{\mathtt{true}}$ Peak bin probability (PBP) of state-binned BS
$\mu_{\mathtt{net}}$ Estimated PBP from all validated samples on the network
$\epsilon$ Desired accuracy of the estimated PBP wrt. $\mu_{\mathtt{true}}$
@@
$N^{(\mathtt{sb})}_{tot}$ The number of samples needed across the network such that $|\mu_{\mathtt{net}}-\mu_{\mathtt{true}}|\leq\epsilon$ with high confidence $N^{(\mathtt{sb})}_{\rm tot}=1.8\times 10^{5}{d^{(\mathtt{sb})}}^{7/2}$
$\gamma$ $100(1-\gamma)\%$ is the confidence interval for $\mu_{\mathtt{true}}$ We assume $\gamma=10^{-4}$
Mode-binned BS $d^{(\mathtt{mb})}$ Number of bins in the mode Hilbert space
$P^{({\mathtt{m}b})}[i]$ Mode-binned probability distribution for miner $i$
@@
$m$ Number of random samples needed to estimate matrix permanent up to an $\delta$ additive estimate with probability at least $p$, using Gurvits’ algorithm $m=\frac{2}{\delta^{2}}\ln(2/(1-p))$
@@
$\beta$ Accuracy parameter used to invalidate a miner’s mode binned distribution based on total variation distance from an estimated mode binned distribution $\widehat{P^{({\mathtt{mb}})}}$ calculated using Gurvits’ algorithm $\mathcal{D}^{({\rm tv})}(\widehat{P^{({\mathtt{mb}})}},P^{({\mathtt{mb}})}[i])\geq 2\beta$
@@
$N^{(\mathtt{mb})}_{tot}$ The number of samples sufficient for a quantum boson sampler to pass the validation test $N^{(\mathtt{mb})}_{{\rm tot}}=2^{14}\frac{\sqrt{\binom{N+d^{(\mathtt{mb})}-1}{N}}}{\beta^{2}}$
$m_{j}[i]$ Number of photons in the $j^{th}$ bin for the $i^{th}$ sample set $s_{i}$
@@
$s_{i}^{(v)}$ Set of samples committed by the $i^{th}$ node in the blockchain network. The $v$ superscript, if any, denotes the samples are validated
$W^{(v)}$ The combined set of samples committed by all the nodes in the blockchain network $W^{(v)}=\bigcup_{i}s^{(v)}_{i}$.
@@
Experimental resources $\eta_{f}$ Combined parameter for single photon generation, coupling, and detection efficiency
$\eta_{t}$ Transmission probability of a single beam-splitter
$R_{0}$ Single photon repetition rate
$R_{q}$ Repetition rate of a quantum boson sampler $R_{q}=(\eta_{f}\eta^{M}_{t})^{N}R_{0}/Ne$
Payoff mechanism $\mu_{i}$ Peak bin probability over the $i^{th}$ verified sample set $s_{i}^{v}$
$t_{i}$ Time when the $i^{th}$ user commits their samples to the network
@@
$p_{i}^{\mathtt{honest}}$ Probability of an honest user (either classical or quantum sampler) to pass the mode-binned BS test
@@
$p_{i}^{\mathtt{cheat}}$ Probability of a user providing non-BS samples to pass the mode-binned BS test
$u_{i}^{\mathtt{honest}}$ Utility of an honest node in the network
$u_{i}^{\mathtt{cheat}}$ Utility of a cheating node in the network
$u_{i}^{\mathtt{nothing}}$ Utility of doing nothing, i.e. not committing any samples
$k^{\mathtt{classical}}$ Cost per honest sample using classical algorithms
@@
$k^{\mathtt{quantum}}\equiv k$ Cost per honest sample using actual BS implementation
$C_{i}$ The cost incurred by the $i^{th}$ node in committing $|s_{i}|$ samples
$P_{i}$ Penalty implemented on the $i^{th}$ node in the network
$R_{i}$ Reward awarded to the $i^{th}$ node in the network